OpenStudy (anonymous):

Evaluate the expression, using the given value of the variable. 3 – 7x + 2x when x = 3 A. –24 B. –12 C. –30 D. –18

OpenStudy (anonymous):

3- 7(3) + 2 (3) = 3-21+6 = -12 It's letter B.